OpenStudy (anonymous):

Please Help!!! I will give medals and will fan!!!! Which of the following methods would be the easiest to use to solve 2x2 + 4x – 3 = 0? A. factoring B. isolating the x2 term and finding the square root of both sides C. using the quadratic formula D. all three methods would be easy and effective

OpenStudy (anonymous):

c or a

OpenStudy (anonymous):

it depand in which method you are better

OpenStudy (anonymous):

I believe the answer is: C) using the quadratic formula.

OpenStudy (anonymous):

can you help me with more plz?

OpenStudy (anonymous):

yes i will try

OpenStudy (anonymous):

Which of the following methods would be the easiest to use to solve x2 + 5x – 6 = 0? A. all three methods would be easy and effective B. using the quadratic formula C. isolating the x2 term and finding the square root of both sides D. factoring

OpenStudy (anonymous):

C. isolating the x2 term and finding the square root of both sides

OpenStudy (anonymous):

d it is easy to factirise and it may take time to solve by formula

OpenStudy (anonymous):

ummmm not sure which one

OpenStudy (anonymous):

Solve. x2 + 4x = 0 A. {0, –4} B. {–4, 0, 4} C. {0, –16} D. {0, 4}

OpenStudy (anonymous):

i think A.

OpenStudy (anonymous):

a

OpenStudy (anonymous):

Yep, A.

OpenStudy (anonymous):

Louis is planning to mow a rectangular yard that is 20 feet longer than it wide. He wants to find its width if its area is 1125 square feet. Which equation models this problem? A. 20w2 = 1125 B. w2 + 20w = 1125 C. w2 – 20w = 1125 D. w + 20w = 1125 Last one

OpenStudy (anonymous):

b

OpenStudy (anonymous):

why

OpenStudy (anonymous):

@Moerdred is he right

OpenStudy (anonymous):

i will explain

OpenStudy (anonymous):

let w is width ok

OpenStudy (anonymous):

so length is 20 + w

OpenStudy (anonymous):

1125 = ( l *b )

OpenStudy (anonymous):

1125 = ( 20 + b ) ( b)

OpenStudy (anonymous):

1125 = 20 w + w^ 2 do you get

OpenStudy (anonymous):

yes thank you

OpenStudy (anonymous):

no problem